OpenStudy (anonymous):

Two angles are complementary if the sum of their measures is 90 degrees. Find two complementary angles such that one of the angles is 141 degrees less than 2 times the other angle. (Round to two decimal places if necessary.)

OpenStudy (anonymous):

Let one angle be x Other angle = 2x-141 And we know, x + (2x - 141) = 90 3x = 231 x = 77 y= 90-77 = 13